Need help:
Use the union rule to answer the question.
If n(A) = 5, n(B) = 11, and n(A ∩ B) = 3; what is n(A u B)?
A) 14
B) 16
C) 13
D) 12

Answer by user_dude2008(1862)

n(A u B) = n(A) + n(B) - n(A ∩ B)

n(A u B) = 5 + 11 - 3

n(A u B) = 16 - 3

n(A u B) = 13

Ans: C) 13